OpenStudy (anonymous):

The sum of two positive numbers is 10. Find the numbers if their squares is a minimum

OpenStudy (anonymous):

Here's one way to do it, but it's not what your looking for. Let x and y be a positive even integer. Then x = 2k for some integer k, y= 2L for some integer L. x + y = 10 (2k) + (2L) = 10 2(k+L) = 10 exp: 2(3+2) = 10

OpenStudy (anonymous):

Thanks StarJ. Although I don't think this is what I'm looking for. currently doing optimisation in year 12 maths B. I know x + y = 10, therefore y = 10 - x but not sure where to go from there.. thanks anyway :)

OpenStudy (noelgreco):

It's not clear to me what "...if their squares is a minimum" means.

OpenStudy (asnaseer):

@siklad96 - you have started correctly by stating that:\[x+y=10\]\[\therefore y=10-x\]Now the question states that you must minimize the sum of their squares. So lets call the sum of the squares some function \(f\) given by:\[f=x^2+y^2\]The aim is to minimize \(f\) - understand so far?

OpenStudy (anonymous):

yes, I understand that because there are two squares therefore x and y are squared, please continue :)

OpenStudy (asnaseer):

good. Now the first step is to reduce the number of variables. From the first condition given to you, we know that:\[y=10-x\]We can therefore substitute this into the expression for \(f\) to obtain:\[f=x^2+y^2=x^2+(10-x)^2=x^2+100-20x+x^2=2x^2-20x+100\] Now, do you know how to find the minimum value for \(f\)?

OpenStudy (anonymous):

I'm pretty sure its the quadratic formula? or you can factorise? correct me if I'm wrong, and how did you end up with \[x^2 + 100 - 20x\]?

OpenStudy (asnaseer):

\[(10-x)^2=(10-x)(10-x)=100-10x-10x+x^2=100-20x+x^2\]

OpenStudy (asnaseer):

so we end up with a quadratic equation (as you correctly recognised):\[f=2x^2-20x+100\]Do you know how to find the minimum value for an equation like this?

OpenStudy (asnaseer):

What optimisation techniques have you been taught?

OpenStudy (anonymous):

\[-b \sqrt{b^2 + 4ac}/2a \] ?

OpenStudy (asnaseer):

no

OpenStudy (asnaseer):

you are not being asked to solve a quadratic equation by finding its roots. If that was the case you would have something like:\[2x^2-20x+100=0\]and you would be asked to find the roots of this equation. This is where you can either factor or use the quadratic equation.

OpenStudy (anonymous):

find the derivative?

OpenStudy (asnaseer):

What you have in this question is a function \(f\) defined as:\[f=2x^2-20x+100\]and you are being asked to find the value of \(x\) that will minimize this function.

OpenStudy (asnaseer):

yes - now you are thinking along the right lines

OpenStudy (anonymous):

4x - 20 ?

OpenStudy (anonymous):

so x = 5

OpenStudy (asnaseer):

what should the derivative be equal to in order to have a minimum (or maximum) value?

OpenStudy (anonymous):

would it be 10

OpenStudy (asnaseer):